If I have 1 mole of CO2 how many molecules of CO2 do I have?
1
Expert's answer
2020-03-17T09:19:22-0400

The number of molecules of CO2 is:

"N=n*N\tiny A"

"N= 1 mol * 6,02 * 10^{23} mol^{-1} = 6,02 * 10^{23}"

The answer is "6,02 * 10^{23}" molecules